OpenStudy (anonymous):

Please factor 6m^2-mn-12n^2, can someone give a hand to this question

OpenStudy (luigi0210):

Find common factors, do you see any?

OpenStudy (anonymous):

\[6m^2-mn-12n^2 = 6m^2-9mn+8mn-12n^2=3m(2m-3n)+4n(2m-3n)\] \[=(3m+4n)(2m-3n)\] Thus (3m+4n) and (2m-3n) are the required factors of the given expression.